Will the volume of a gas increase, decrease, or remain unchanged for the following set of changes? The pressure is decreased from 6.33 atm to 1.05 atm, while the temperature is increased from 123°C to 234°C.


a) The volume will increase

b) The volume will decrease

c) The volume will not change


1
Expert's answer
2021-03-02T04:55:05-0500

PV/T = P0V0/T0

1.05*V/234=6.33*V0/123

0.0045*V=0.0515*V0

V/V0=0.0515/0.0045

V/V0=11.4.

So, the answer is a) the volume will increase